OpenStudy (imnotgoodatmath):

What is the x-intercept of the line given by the equation x - 4y = 20? The x-intercept is 0. The x-intercept is –4. The x-intercept is 5. The x-intercept is 20.

OpenStudy (zale101):

When finding the x-intercept you replace y by zero and solve for x. In your equation, you first need to set the equation in the form of slope intercept y=mx+b so x can be easier to solve for.

OpenStudy (anonymous):

no you don't just erase the \(y\) and get \(x=20\)

OpenStudy (anonymous):

n order to find the x-intercept, you must substitute the y as 0, as any x-intercept has a y-value of 0. Therefore... x - 4y = 20 x - 4(0) = 20 x = 20 The x-intercept is 20.

OpenStudy (anonymous):

did this help @ImNotGoodAtMath

OpenStudy (anonymous):

did this help at all